jun 18, 1999 owl shift summary

Richard Kessler (rkessler@fnal.gov)
Fri, 18 Jun 1999 13:28:53 -0500 (CDT)

Shift Summary for OWL shift, June 18, 1999

On Shift:RK PT

Hours of beam delivered at all intensities: 7 (note: 7 hr shift)
Hours of beam delivered at nominal intensity: 7
Hours of E832 data (special runs not included): 6.7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 15 minutes to change tapes

========================================================================
Runs ended this shift:
Run 13643 - 9.5M events - BEAM_832 (15K vac00)

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth E832 data taking.

We finally found that beam arrives 150 msec BEFORE the T5
signal! This was found by triggering the scope on T5 and then
looking at a REG*V coincidence. If MCR cannot fix the T5 signal
we will inhibit triggers betweeen T4 and T5.

L3_drivers re-made between runs with DC monitor fixes
(see run 13644).

2 ns was added to the ETOT L1 trigger source (before run 13644).

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Rad safety crew expected during day shift.